OpenStudy (anonymous):

what is the solution to the rational expression... (5/x^2+3x+2)+(1/x+2)=(1/3x+3)... is it 9, 5/2, -7, -8

OpenStudy (callisto):

\[\frac{5}{x^2+3x+2}+\frac{1}{x+2}=\frac{1}{3x+3}\]\[\frac{5}{(x+2)(x+1)}+\frac{1}{x+2}=\frac{1}{3(x+1)}\]\[\frac{5}{(x+2)(x+1)}+\frac{1(x+1)}{(x+2)(x+1)}=\frac{1}{3(x+1)}\]\[\frac{5+(x+1)}{(x+2)}=\frac{1}{3}\]\[\frac{(x+6)}{(x+2)}=\frac{1}{3}\]3(x+6) = x+2 3x+18 = x+2 2x = -16 x=?

OpenStudy (anonymous):

-8